OpenStudy (anonymous):

Determine if the lines are parallel, perpendicular, same line, or none of the above. y=-3/5x+2and 3x+5y=20

OpenStudy (anonymous):

find the slope of both equations slope of first line is m1=-3/5 slope of second line is m2=-3/5 since slopes are same so the lines are parallel